# Active Learning of Deep Neural Networks via Gradient-Free Cutting Planes

### v1

@@ 10 Acknowledgments
Appendix A Key Definitions and Deferred Theorems
A.1 Key Definition
@@
As the readers may see in the definition, $P$, the number of regions formed by the hyperplane arrangement, increases with both the dimension and the size of the dataset $X$. To reduce computational costs of our cutting-plane AL algorithm, as we will discuss in Appendix F.1, we sample the number of regions in the partition instead of using the entire $\mathcal{S}$.
@@ A.2 Key Theorems
Appendix B Deferred Algorithms
B.1 More on Cutting-Plane AL for Binary Classification
B.1.1 Linear Model
@@
B.1.2 NN Model with Limited Queries
@@
B.1.3 NN Model with Inexact Cut
Algorithm 5 summarizes our proposed cutting-plane active learning algorithm under the third setup metioned in Section 7.1. Under this scenario, the cutting-plane oracle has access to limited queries. However, in contrast to Algorithm 4, this cutting-plane AL algorithm always performs the cut regardless of whether the queried center mis-classifies the data point returned by the oracle. This is an interesting extension to consider. On the one hand, it can possibly speed up Algorithm 4 in making a decision boundary as the cut is effective in every iteration. On the other hand, however, the algorithm’s lack of discern for the correctness of the queried candidate presents a non-trivial challenge to evaluate its convergence rate and whether it still maintains convergence guarantees. It turns out that we can still ensure convergence in the case of Algorithm 5, and the convergence rate can be quantified by measuring the “inexactness” of the cut in relation to a cut which directly passes through the queried center. We refer the readers to a detailed discussion on this matter in Appendix E.2. We would also like to emphasize that Algorithm 2, 5, and 4, although written for binary classification, can be easily extended to multi-class by using, for instance, the “one-versus-all” strategy. See Appendix C.2 for details.
While it may be intuitive for the optimal performance of the cutting-plane AL algorithms to translate from binary classification to the multi-class case, it is not entirely evident for us to expect similar performance of the algorithms for regression tasks, where the number of classes $K\rightarrow\infty$. What is surprising is that our cutting-plane AL still maintains its optimal performance on regression tasks, as evidenced by the synthetic toy example using quadratic regression in Section 8. Nevertheless, one can argue that the result is not so surprising after all as it is to be expected in theory due to intuition explained in Appendix C.2.
@@ B.1.3 NN Model with Inexact Cut
B.2 More on Cutting-Plane AL for Regression
B.2.1 NN Model with Limited Queries
@@ B.2.2 Linear Model
B.3 Minimal Margin Query Strategy
@@
Appendix C Key Generalization to Cutting-Plane AL
@@
C.1 Relaxed Data Distribution Requirement
@@
The above result can be extended to any $f\in C([-R,R]^{d})$, see Theorem 2.5 in Shen et al. (2022) for discussion. Therefore, for any fixed dimension $d$, as we increase number of hidden neurons, we are guaranteed to approximate $f$ a.e. (under proper condition on $\omega_{f}$). Thus the assumption on training data can be relaxed to existence of some $\{W_{1},b_{1},W_{2},b_{2}\}$ such that ${\mathrm{sign}}(\phi(x))=y$, which happens almost surely. A minor discrepancy here is that the $\phi$ being considered in above Theorem incorporates the bias terms $b_{1},b_{2}$ while our Theorem 5.1 considers two-layer NN of form $(xW_{1})_{+}W_{2}$. We note here that $(xW_{1})_{+}W_{2}$ is essentially the same as the one with layer-wise bias. To see this, we can append the data $x$ by a $1$-value entry to accommodate for $b_{1}$. After this modification, we can always separate a neuron of form $(XW_{1i})_{+}W_{2i}$, set $W_{1(d+1)}=1$ and zeros elsewhere, and $W_{2i}$ would then accommodate for bias $b_{2}.$ Therefore, our relaxed requirement on training data still persists. For deeper ReLU NNs, similar uniform approximation capability has been established priorly and furthermore the general NN form we considered in Theorem 5.2 incorporates the form with layer-wise bias and thus we still have such relaxation on training data distribution, see Appendix D.3 for more explanation.
C.2 From Classification to Regression
@@
Appendix D Deferred Proofs and Extensions in Section 5
D.1 Proof of Theorem 5.1
@@ Appendix D Deferred Proofs and Extensions in Section 5
D.3 Explanation of Bias Term for General Case
@@
D.4 Activation Pattern Subsampling and Iterative Filtering.
Here we detail more about our hyperplane selection scheme. Take two-layer ReLU model for example, in order to find the activation pattern $\mathcal{D}$ corresponding to the hidden layer, one needs to exhaust the set $\{\text{diag}({\mathbbm{1}}\{Xu\geq 0\})\}$ for all $u\in\mathbb{R}^{d}.$ In our experiments, we adopt a heuristic subsampling procedure, i.e., we usually set a moderate number of hidden neurons $m_{1}$, and we sample a set of Gaussian random vectors $\{u_{1},u_{2},\cdots,u_{n_{1}}\}$ with some random $n_{1}>m_{1}.$ Then we take the set $\{\text{diag}({\mathbbm{1}}\{Xu_{i}\geq 0\})|i\in[n_{1}]\}.$ If |$|\{\text{diag}({\mathbbm{1}}\{Xu_{i}\geq 0\})|i\in[n_{1}]\}|>m_{1},$ we take a subset of $m_{1}$ activation patterns there, if not, we increase $n_{1}$ and redo all prior steps until we hit some satisfactory $n_{1}$. This heuristic method always works well in our experiments, see Appendix F.1 for more about our implementation details.
@@
Appendix E Deferred Proofs and Extensions in Section 7
E.1 Deferred Proof of Theorem 7.1
@@ Appendix E Deferred Proofs and Extensions in Section 7
E.2 Extension to Algorithm 4: The Case of Inexact Cuts and Convergence
@@
Theorem E.1 (Convergence with Inexact Cuts of Center of Gravity).
@@
To justify that the derived ratio $\textbf{vol}(\mathcal{T}_{1})/\textbf{vol}(\mathcal{T})$ is valid, i.e. $\textbf{vol}(\mathcal{T}_{1})/\textbf{vol}(\mathcal{T})\in[0,1]$, it suffices to show that $e^{-1}(1-\tilde{\lambda})^{d}\in[0,1]$, or $0<\tilde{\lambda}\leq 1$.
Define coefficient
@@
Given Theorem E.1, at each step $t$ in Algorithm 5, at least a volume of magnitude $e^{-1}(1-\tilde{\lambda})^{d}\textbf{vol}(\mathcal{T}^{t})$ is cut away. The volume of $\mathcal{T}^{t}$ after $t$ iterations is bounded by:
@@
Theorem E.1 allows us to quantify the volume reduction rate as a function of the inexactness of cuts through the Euclidean distance between the centroid and the cutting hyperplane, which is monitored by the query sampling method. In particular, in the case of inexact cuts in Algorithm 5 with a minimal margin query sampling scheme, the volume reduction rate converges asymptotically in the number of training data supplied to the cutting-plane oracle to the rate in Proposition 1. This result is stated in the following corollary.
@@
By Theorem E.1, given an inexact polyhedron cut in Algorithm 5 distance $h$ away from the centroid, the volume of the remaining set is upper bounded by the following:
@@
E.3 Convergence w.r.t. Center of the Maximum Volume Ellipsoid
@@
Theorem E.2 (Convergence with Center of MVE).
@@
Appendix F Experiments Supplementals
F.1 Implementation Details
In this section, we provide an overview of the implementation details. For specific aspects, such as baseline implementation and the cutting-plane AL method for regression, we direct readers to the corresponding sections in Appendix F.
In our experiments, we follow exactly the same algorithm workflow as in Algorithm 1 for our model training and Algorithm 4 for active learning with limited queries, which is the version we used conducting all of our experiments in Section 8. In the training of Algorithm 4, we implement the “center” function for analytic center retrieval due to its simple computation formula (see Definition 1). Since the center retrieval problem is of convex minimization form, we solve it with CVXPY (Diamond & Boyd, 2016) and default to MOSEK (MOSEK ApS, 2024) as our solver. In our experiments, MOSEK is able to handle all encountered convex optimization programs, given that they are feasible. We involve all training data cuts while we note that some cut dropping method may help alleviate computation overhead, i.e., one may keep the last several cuts only, see Section 2.5 in (Parshakova et al., 2023) for a discussion on constraint dropping. For activation pattern generation, for two-layer model experiment, say we want $P=m$ patterns, we always generate a set of standard Gaussian vectors $\{u_{i},i\in[n]\}$ for some $n>m$, compute the induced pattern set $\{\mathcal{D}_{i}=\text{diag}({\mathbbm{1}}\{Xu_{i}\geq 0\})\}$, and take $m$ non-duplicate patterns out of it. Furthermore, in our implementation of Algorithm 4, we included a few additional checks for computational efficiency, like skipping the centering step when no cut has been performed before, or book-keeping of discarded data points that should not be re-considered immediately. For IMDB experiments, we randomly pick $50$ training data points and $20$ test data points for ease of computation, and we use minimal margin query strategy described in Appendix B.3. Through all our data experiments, random seed is fixed to be 00 except for error bar plots, where we always take random seeds $\{0,1,2,3,4\}$.
F.2 Final Solve Regularization in Cutting-plane AL
@@
$\begin{split}p^{*}=&\min_{\{u_{i},{u^{\prime}}_{i}\}_{i=1}^{P},u_{i},{u^{\prime}}_{i}\in\mathbb{R}^{d}\forall i}\frac{1}{2}\|\sum_{i=1}^{P}D(S_{i})X({u^{\prime}}_{i}-u_{i})-y\|^{2}_{2}\\ &+\beta(\sum_{i=1}^{P}(\|u_{i}\|_{2}+\|u^{\prime}_{i}\|_{2}))\end{split}$ (27)
@@
First, using the same spiral dataset of randomly selected $80$ points from $100$ points Spiral $(k_{1}=13,k_{2}=0.5,n_{\text{shape}}=50)$ (detailed in Appendix F.4) and $\beta=0.001$, the convex solve achieves an objective value of 0.0008, while stochastic gradient descent levels out at an objective value greater than 0.001. The final decision boundary on the spiral using the convex solve (left) and SGD (right) is shown in Figure 5, where we have used marker “x” to indicate the train points passed into each solver and triangle marker to indicate test points.
@@
$\sum_{i=1}^{P}D(S_{i})X({u^{\prime}}_{i}-u_{i}):=f^{\text{two-layer}}(X;\theta)=[X^{1}-X^{1}\ ...\ X^{P}-X^{P}]\theta,$
@@
$\begin{split}f^{\text{obj}}(\mathcal{D};\theta,\beta)=&\min_{\theta}\frac{1}{2}\|\sum_{i=1}^{P}(D(S_{i})X)_{\mathcal{D}}({u^{\prime}}_{i}-u_{i})-y_{\mathcal{D}}\|^{2}_{2}\\ &+\beta(\sum_{i=1}^{P}(\|u_{i}\|_{2}+\|u^{\prime}_{i}\|_{2})),\end{split}$ (28)
@@
Upon examining Algorithm 8 and the objective function of the equivalent convex program $f^{\text{obj}}(\mathcal{D}_{\text{AL}};\theta^{t},\beta)$ in 27, a key difference is the introduction of a regularization term, i.e. $\beta(\sum_{i=1}^{P}(\|u_{i}\|_{2}+\|u^{\prime}_{i}\|_{2}))$. In more complex tasks, the inclusion of the final convex solve, and thus an additional regularization, tends to aid the cutting-plane AL method in achieving faster convergence in the number of queries. This is the case, for example, for the spiral dataset in our experiment (see Appendix F.4). However, for simpler tasks, such as the quadratic regression prediction, adding regularization could slow down the convergence (see Appendix F.5). Therefore, in our experiments, we emphasize that we have used the best cutting-plane AL method, chosen from the version with or without the final convex solve. This is explicitly noted in the supplementary sections, namely Appendix F.4 for the spiral experiment and Appendix F.5 for the quadratic regression task. We also remark that due to the limitation in theory on the convexification of neural networks, such a convex solve is only viable up to a three-layer ReLU network.
F.3 Active Learning Baselines

### v3

@@ 9 Acknowledgments
Appendix A Prior Work
Cutting-Plane Method.
Active Learning.
With the fast growth of current model sizes, a large amounts of data is need for training an effective deep networks. Compared to just feeding the model with a set of randomly-selected training samples, how to select the most informative data for each training iteration becomes a critical question, especially in neural models (Zhu & Nowak, 2022; Karzand & Nowak, 2020). Efficient data sampling is thus increasingly important, especially for current RL-based language model training schemes. With the increasing power of current pretrained models, researchers observed that since most of simple questions can be addressed correctly by the model already, these questions are less important for further boosting the language models’ capacity. Therefore, actively selecting questions that LLM cannot address correctly is key to current LLM training (Lightman et al., 2023). Based on different learning settings, active learning strategies can be divided into: stream-based selective sampling used when the data is generated continuously (Woodward & Finn, 2017); pool-based sampling used when a pool of unlabeled data is presented (Gal et al., 2017b); query synthesis methods used when new samples can be generated for labeling. Based on different information measuring schemes, active learning algorithms can be further divided into uncertainty sampling (Lewis & Gale, 1994b) which selects data samples to reduce prediction uncertainty, query-by-committee sampling which involves multiple models for data selection, diversity-weighted method which selects the most diverse data sample, and expected error reduction method (Mussmann et al., 2022) which selects samples to best reduce models’ expected prediction error.
Convex NN.
Appendix B Key Definitions and Deferred Theorems
B.1 Key Definition
@@
As the readers may see in the definition, $P$, the number of regions formed by the hyperplane arrangement, increases with both the dimension and the size of the dataset $X$. To reduce computational costs of our cutting-plane AL algorithm, as we will discuss in Appendix G.1, we sample the number of regions in the partition instead of using the entire $\mathcal{S}$.
@@ B.2 Key Theorems
Appendix C Deferred Algorithms
C.1 More on Cutting-Plane AL for Binary Classification
C.1.1 Linear Model
@@
C.1.2 NN Model with Limited Queries
@@
C.1.3 NN Model with Inexact Cut
Algorithm 5 summarizes our proposed cutting-plane active learning algorithm under the third setup metioned in Section 6.1. Under this scenario, the cutting-plane oracle has access to limited queries. However, in contrast to Algorithm 4, this cutting-plane AL algorithm always performs the cut regardless of whether the queried center mis-classifies the data point returned by the oracle. This is an interesting extension to consider. On the one hand, it can possibly speed up Algorithm 4 in making a decision boundary as the cut is effective in every iteration. On the other hand, however, the algorithm’s lack of discern for the correctness of the queried candidate presents a non-trivial challenge to evaluate its convergence rate and whether it still maintains convergence guarantees. It turns out that we can still ensure convergence in the case of Algorithm 5, and the convergence rate can be quantified by measuring the “inexactness” of the cut in relation to a cut which directly passes through the queried center. We refer the readers to a detailed discussion on this matter in Appendix F.2. We would also like to emphasize that Algorithm 2, 5, and 4, although written for binary classification, can be easily extended to multi-class by using, for instance, the “one-versus-all” strategy. See Appendix D.2 for details.
While it may be intuitive for the optimal performance of the cutting-plane AL algorithms to translate from binary classification to the multi-class case, it is not entirely evident for us to expect similar performance of the algorithms for regression tasks, where the number of classes $K\rightarrow\infty$. What is surprising is that our cutting-plane AL still maintains its optimal performance on regression tasks, as evidenced by the synthetic toy example using quadratic regression in Section 7. Nevertheless, one can argue that the result is not so surprising after all as it is to be expected in theory due to intuition explained in Appendix D.2.
@@ C.1.3 NN Model with Inexact Cut
C.2 More on Cutting-Plane AL for Regression
C.2.1 NN Model with Limited Queries
@@ C.2.2 Linear Model
C.3 Minimal Margin Query Strategy
@@
Appendix D Key Generalization to Cutting-Plane AL
@@
D.1 Relaxed Data Distribution Requirement
@@
The above result can be extended to any $f\in C([-R,R]^{d})$, see Theorem 2.5 in Shen et al. (2022) for discussion. Therefore, for any fixed dimension $d$, as we increase number of hidden neurons, we are guaranteed to approximate $f$ a.e. (under proper condition on $\omega_{f}$). Thus the assumption on training data can be relaxed to existence of some $\{W_{1},b_{1},W_{2},b_{2}\}$ such that ${\mathrm{sign}}(\phi(x))=y$, which happens almost surely. A minor discrepancy here is that the $\phi$ being considered in above Theorem incorporates the bias terms $b_{1},b_{2}$ while our Theorem 4.1 considers two-layer NN of form $(xW_{1})_{+}W_{2}$. We note here that $(xW_{1})_{+}W_{2}$ is essentially the same as the one with layer-wise bias. To see this, we can append the data $x$ by a $1$-value entry to accommodate for $b_{1}$. After this modification, we can always separate a neuron of form $(XW_{1i})_{+}W_{2i}$, set $W_{1(d+1)}=1$ and zeros elsewhere, and $W_{2i}$ would then accommodate for bias $b_{2}.$ Therefore, our relaxed requirement on training data still persists. For deeper ReLU NNs, similar uniform approximation capability has been established priorly and furthermore the general NN form we considered in Theorem 4.2 incorporates the form with layer-wise bias and thus we still have such relaxation on training data distribution, see Appendix E.3 for more explanation.
D.2 From Classification to Regression
@@
Appendix E Deferred Proofs and Extensions in Section 4
E.1 Proof of Theorem 4.1
@@ Appendix E Deferred Proofs and Extensions in Section 4
E.3 Explanation of Bias Term for General Case
@@
E.4 Activation Pattern Subsampling and Iterative Filtering.
Here we detail more about our hyperplane selection scheme. Take two-layer ReLU model for example, in order to find the activation pattern $\mathcal{D}$ corresponding to the hidden layer, one needs to exhaust the set $\{\text{diag}({\mathbbm{1}}\{Xu\geq 0\})\}$ for all $u\in\mathbb{R}^{d}.$ In our experiments, we adopt a heuristic subsampling procedure, i.e., we usually set a moderate number of hidden neurons $m_{1}$, and we sample a set of Gaussian random vectors $\{u_{1},u_{2},\cdots,u_{n_{1}}\}$ with some random $n_{1}>m_{1}.$ Then we take the set $\{\text{diag}({\mathbbm{1}}\{Xu_{i}\geq 0\})|i\in[n_{1}]\}.$ If |$|\{\text{diag}({\mathbbm{1}}\{Xu_{i}\geq 0\})|i\in[n_{1}]\}|>m_{1},$ we take a subset of $m_{1}$ activation patterns there, if not, we increase $n_{1}$ and redo all prior steps until we hit some satisfactory $n_{1}$. This heuristic method always works well in our experiments, see Appendix G.1 for more about our implementation details.
@@
Appendix F Deferred Proofs and Extensions in Section 6
F.1 Deferred Proof of Theorem 6.1
@@ Appendix F Deferred Proofs and Extensions in Section 6
F.2 Extension to Algorithm 4: The Case of Inexact Cuts and Convergence
@@
Theorem F.1 (Convergence with Inexact Cuts of Center of Gravity).
@@
To justify that the derived ratio $\textbf{vol}(\mathcal{T}_{1})/\textbf{vol}(\mathcal{T})$ is valid, i.e. $\textbf{vol}(\mathcal{T}_{1})/\textbf{vol}(\mathcal{T})\in[0,1]$, it suffices to show that $e^{-1}(1-\tilde{\lambda})^{d}\in[0,1]$ for all given $d>0$, which necessarily holds if $0<\tilde{\lambda}\leq 1$.
To see that $0<\tilde{\lambda}\leq 1$, we first define coefficient
@@
Given Theorem F.1, at each step $t$ in Algorithm 5, at least a volume of magnitude $e^{-1}(1-\tilde{\lambda})^{d}\textbf{vol}(\mathcal{T}^{t})$ is cut away. The volume of $\mathcal{T}^{t}$ after $t$ iterations is bounded by:
@@
Theorem F.1 allows us to quantify the volume reduction rate as a function of the inexactness of cuts through the Euclidean distance between the centroid and the cutting hyperplane, which is monitored by the query sampling method. In particular, in the case of inexact cuts in Algorithm 5 with a minimal margin query sampling scheme, the volume reduction rate converges asymptotically in the number of training data supplied to the cutting-plane oracle to the rate in Proposition 1. This result is stated in the following corollary.
@@
By Theorem F.1, given an inexact polyhedron cut in Algorithm 5 distance $h$ away from the centroid, the volume of the remaining set is upper bounded by the following:
@@
F.3 Convergence w.r.t. Center of the Maximum Volume Ellipsoid
@@
Theorem F.2 (Convergence with Center of MVE).
@@
Appendix G Experiments Supplementals
G.1 Implementation Details
In this section, we provide an overview of the implementation details. For specific aspects, such as baseline implementation and the cutting-plane AL method for regression, we direct readers to the corresponding sections in Appendix G.
In our experiments, we follow exactly the same algorithm workflow as in Algorithm 1 for our model training and Algorithm 4 for active learning with limited queries, which is the version we used conducting all of our experiments in Section 7. In the training of Algorithm 4, we implement the “center” function for analytic center retrieval due to its simple computation formula (see Definition 1). Since the center retrieval problem is of convex minimization form, we solve it with CVXPY (Diamond & Boyd, 2016) and default to MOSEK (MOSEK ApS, 2024) as our solver. In our experiments, MOSEK is able to handle all encountered convex optimization programs, given that they are feasible. We involve all training data cuts while we note that some cut dropping method may help alleviate computation overhead, i.e., one may keep the last several cuts only, see Section 2.5 in (Parshakova et al., 2023) for a discussion on constraint dropping. For activation pattern generation, for two-layer model experiment, say we want $P=m$ patterns, we always generate a set of standard Gaussian vectors $\{u_{i},i\in[n]\}$ for some $n>m$, compute the induced pattern set $\{\mathcal{D}_{i}=\text{diag}({\mathbbm{1}}\{Xu_{i}\geq 0\})\}$, and take $m$ non-duplicate patterns out of it. Furthermore, in our implementation of Algorithm 4, we included a few additional checks for computational efficiency, like skipping the centering step when no cut has been performed before, or book-keeping of discarded data points that should not be re-considered immediately. For IMDB experiments, we randomly pick $50$ training data points and $20$ test data points for ease of computation, and we use minimal margin query strategy described in Appendix C.3. Through all our data experiments, random seed is fixed to be 00 except for error bar plots, where we always take random seeds $\{0,1,2,3,4\}$.
G.2 Final Solve Regularization in Cutting-plane AL
@@
$\begin{split}p^{*}=&\min_{\{u_{i},{u^{\prime}}_{i}\}_{i=1}^{P},u_{i},{u^{\prime}}_{i}\in\mathbb{R}^{d}\forall i}\frac{1}{2}\|\sum_{i=1}^{P}D_{i}X({u^{\prime}}_{i}-u_{i})-y\|^{2}_{2}\\ &+\beta(\sum_{i=1}^{P}(\|u_{i}\|_{2}+\|u^{\prime}_{i}\|_{2}))\end{split}$ (27)
@@
First, using the same spiral dataset of randomly selected $80$ points from $100$ points Spiral $(k_{1}=13,k_{2}=0.5,n_{\text{shape}}=50)$ (detailed in Appendix G.4) and $\beta=0.001$, the convex solve achieves an objective value of 0.0008, while stochastic gradient descent levels out at an objective value greater than 0.001. The final decision boundary on the spiral using the convex solve (left) and SGD (right) is shown in Figure 5, where we have used marker “x” to indicate the train points passed into each solver and triangle marker to indicate test points.
@@
$\sum_{i=1}^{P}D_{i}X({u^{\prime}}_{i}-u_{i}):=f^{\text{two-layer}}(X;\theta)=[X^{1}-X^{1}\ ...\ X^{P}-X^{P}]\theta,$
@@
$\begin{split}f^{\text{obj}}(\mathcal{D};\theta,\beta)=&\min_{\theta}\frac{1}{2}\|\sum_{i=1}^{P}(D_{i}X)_{\mathcal{D}}({u^{\prime}}_{i}-u_{i})-y_{\mathcal{D}}\|^{2}_{2}+\beta(\sum_{i=1}^{P}(\|u_{i}\|_{2}+\|u^{\prime}_{i}\|_{2})),\end{split}$ (28)
@@
Upon examining Algorithm 8 and the objective function of the equivalent convex program $f^{\text{obj}}(\mathcal{D}_{\text{AL}};\theta^{t},\beta)$ in 27, a key difference is the introduction of a regularization term, i.e. $\beta(\sum_{i=1}^{P}(\|u_{i}\|_{2}+\|u^{\prime}_{i}\|_{2}))$. In more complex tasks, the inclusion of the final convex solve, and thus an additional regularization, tends to aid the cutting-plane AL method in achieving faster convergence in the number of queries. This is the case, for example, for the spiral dataset in our experiment (see Appendix G.4). However, for simpler tasks, such as the quadratic regression prediction, adding regularization could slow down the convergence (see Appendix G.5). Therefore, in our experiments, we emphasize that we have used the best cutting-plane AL method, chosen from the version with or without the final convex solve. This is explicitly noted in the supplementary sections, namely Appendix G.4 for the spiral experiment and Appendix G.5 for the quadratic regression task. We also remark that due to the limitation in theory on the convexification of neural networks, such a convex solve is only viable up to a three-layer ReLU network.
G.3 Active Learning Baselines